# Groups of matrices with approximately submultiplicative spectra

### v3

@@ 4.2. Remarks on a representation-theoretic version of approximate submultiplicativity and linear bounds
The class ($\hat{\mathrm{s}}$) of groups $G$ with the property that all irreducible representations of all subgroups are submultiplicative was introduced in [6] and then further studied in [3]. In analogy we make the following definition.
@@
Let $\underline{\varepsilon}=(\varepsilon_{n})_{n=1}^{\infty}$ be a sequence of nonnegative real numbers and let $G$ be an abstract group. We say that $G$ is in class $(\underline{\varepsilon}-\widehat{\mathrm{s}})$ if for every $n$, every subgroup $H$ of $G$, and every irreducible representation $\rho\colon H\to\mathcal{M}_{n}(\mathbb{C})$, the image $\rho(H)$ is $\varepsilon_{n}$-submultiplicative. We say that $G$ is in class $(\underline{\varepsilon}-\widehat{\mathrm{as}})$ if for every $n$ and every subgroup $H$ of $G$ we have that the image of every irreducible unitary representation $\rho\colon H\to\mathcal{M}_{n}(\mathbb{C})$ is $\varepsilon_{n}$-ASM.
We can also define analogous notions for any fixed representation: a representation $\rho\colon G\to\mathcal{M}_{n}(\mathbb{C})$ is $\underline{\varepsilon}-\widehat{\mathrm{s}}$, or $\underline{\varepsilon}-\widehat{\mathrm{as}}$ respectively, if for every subgroup $H$ and every irreducible subrepresentation of $H$, $\rho_{W}\colon H\to\operatorname{End}(W)\simeq\mathcal{M}_{m}(\mathbb{C})$, the image $\rho_{W}(H)$ is $\varepsilon_{m}$-submultiplicative, or $\varepsilon_{m}$-ASM respectively (here $W$ is a minimal invariant subspace for $\rho(H)$ of dimension $m$).
@@
In other words, $q\in Q(p)$ precisely when the distance from any fraction $\frac{k}{q}$, $k=1\ldots,q-1$ to the closest fraction of the form $\frac{j}{p}$, $j=0,\ldots,p$ is smaller then $\varepsilon_{p}$. Note that if $q>\frac{1}{2\delta_{p}}$, then $q\not\in Q(p)$. Hence the set $Q(p)$ is always finite.
Assume now that $G$ is a finite minimal nonabelian group whose order is divisible by two primes $p$, $q$ (not necessarily distinct) and that $[G,G]$ is a $q$-group (in the language of [10] we would say that $G$ the $(p,q,f)$-group for some irreducible divisor $f$ of $x^{p}-1\in\mathbb{Z}_{q}[x]$). It is then easy to see (using the same ideas as in the proof of Theorem 6) that $G$ is in class $(\underline{\varepsilon}-\widehat{\mathrm{as}})$ if and only if $q\in Q(p)$. Since $Q(p)$ is finite we can therefore conclude that any compact matrix group (viewed as a representation of itself in the obvious way) satisfying $\underline{\varepsilon}-\widehat{\mathrm{as}}$ must be finite modulo its centre.
We also remark that there are easy examples of sequences $\underline{\varepsilon}$ satisfying the above properties such that every $Q(p)$ also contains primes $q$ different from $p$. In these cases, the corresponding $(p,q,f)$-groups will not be nilpotent. Hence for such $\underline{\varepsilon}$, the class $(\underline{\varepsilon}-\widehat{\mathrm{as}})$ strictly extends the class $(\widehat{\mathrm{s}})$. Hence it is natural to ask the followung question.

### v4

@@ 4.2. Remarks on a representation-theoretic version of approximate submultiplicativity and linear bounds
The class ($\hat{\mathrm{s}}$) of groups $G$ with the property that all irreducible representations of all subgroups are submultiplicative was introduced in [6] and then studied further in [3]. We extend this class of groups in the definition below. Note that in the case of $\varepsilon$-submultiplicativity we are considering all representations, whereas in the case of $\varepsilon$-ASM we restrict ourselves to the class of unitary representations. We remark that for finite groups there is no loss of generality in this restriction, as every representation is equivalent to a unitary representation.
@@
Let $\underline{\varepsilon}=(\varepsilon_{n})_{n=1}^{\infty}$ be a sequence of nonnegative real numbers and let $G$ be an abstract group. We say that $G$ is in class $(\underline{\varepsilon}-\widehat{\mathrm{s}})$ if for every $n$, every subgroup $H$ of $G$, and every irreducible representation $\rho\colon H\to\mathcal{M}_{n}(\mathbb{C})$, the image $\rho(H)$ is $\varepsilon_{n}$-submultiplicative. We say that $G$ is in class $(\underline{\varepsilon}-\widehat{\mathrm{as}})$ if for every $n$ and every subgroup $H$ of $G$, the image of every irreducible unitary representation $\rho\colon H\to\mathcal{M}_{n}(\mathbb{C})$ is $\varepsilon_{n}$-ASM.
We also define the corresponding notions for any fixed representation: a representation $\rho\colon G\to\mathcal{M}_{n}(\mathbb{C})$ is $\underline{\varepsilon}-\widehat{\mathrm{s}}$, or $\underline{\varepsilon}-\widehat{\mathrm{as}}$ respectively (for the latter we additionally assume that $\rho$ is unitary), if for every subgroup $H$ and every irreducible subrepresentation of $H$, $\rho_{W}\colon H\to\operatorname{End}(W)\simeq\mathcal{M}_{m}(\mathbb{C})$, the image $\rho_{W}(H)$ is $\varepsilon_{m}$-submultiplicative, or $\varepsilon_{m}$-ASM, respectively (here $W$ is a minimal invariant subspace for $\rho(H)$ of dimension $m$).
@@
In other words, $q\in Q(p)$ precisely when the distance from any fraction $\frac{k}{q}$, $k=1\ldots,q-1$ to the closest fraction of the form $\frac{j}{p}$, $j=0,\ldots,p$ is smaller than $\varepsilon_{p}$. Note that if $q>\frac{1}{2\delta_{p}}$, then $q\not\in Q(p)$. Therefore, the set $Q(p)$ is always finite.
Assume now that $G$ is a finite minimal nonabelian group whose order is divisible by two primes $p$, $q$ (not necessarily distinct) and that $[G,G]$ is a $q$-group (in the language of [10] we would say that $G$ is the $(p,q,f)$-group for some irreducible divisor $f$ of $x^{p}-1\in\mathbb{Z}_{q}[x]$). It is then easy to see (using the same ideas as in the proof of Theorem 6) that $G$ is in the class $(\underline{\varepsilon}-\widehat{\mathrm{as}})$ if and only if $q\in Q(p)$. Since $Q(p)$ is finite, we can therefore conclude that any compact matrix group (viewed as a representation of itself in the obvious way) satisfying $\underline{\varepsilon}-\widehat{\mathrm{as}}$ must be finite modulo its centre.
We also remark that there are easy examples of sequences $\underline{\varepsilon}$ satisfying the above properties such that every $Q(p)$ also contains primes $q$ different from $p$. In these cases, the corresponding $(p,q,f)$-groups will not be nilpotent. Hence, for such $\underline{\varepsilon}$, the class $(\underline{\varepsilon}-\widehat{\mathrm{as}})$ strictly extends the class $(\widehat{\mathrm{s}})$. Hence, it is natural to ask the following question.
@@
Acknowledgments
We would like to express our gratitude to the referee. Their suggestions significantly improved the paper.
M. Mastnak was supported in part by the NSERC Discovery Grant 371994-2019. Most of the research presented in the paper was done as part of an undergraduate summer research project. The authors thank Saint Mary’s University and NSERC for their generous support.